Write the equation of the line that passes through point (–1, 4) with a slope of 0.
not sure if i did this correct can you help? thnak you

Answer by dolly(163)

Slope of the line = 0
It passes through (-1,4)
Equation of the line is y-y1 = m(x-x1) where m is the slope and (x1,y1) is the point through which it passes.
==> y - 4 = 0(x - (-1))
==> y - 4 = 0
==> y = 4 is the required equation